For what value of c will 3 x^2 - 2x + c = 0 have exactly one distinct real root?
A -2/3 B -1/3 C 0 D 1/3 E 2/3

Answer by mananth(16949)

3x^-2x+c
for c=0
3x^2-2x =0
x(3x-2)=0
3x-2 =0
3x=2
x=2/3